OpenStudy (skyepurple1997):

I'm having so trouble with this problem. I keep getting z-score = -3.9, but the z-score doesn't go that high. Here's the problem: cars are normally distributed Mew=88tons STD=0.9tons B. What is the probability that 50 cars chosen at random will have a mean load weight X bar of less than 87.5 tons of coal? (Round to 4 decimal places.)

OpenStudy (skyepurple1997):

\[\sigma xbar/\sqrt{n} = 0.9/ sq root 50 = 0.127

OpenStudy (skyepurple1997):

p(xbar < 87.5) = 87.5 - 88/0.127 = -3.9

OpenStudy (amistre64):

what are you using to convert z score to probability?

OpenStudy (skyepurple1997):

umm... z = x - mew/ sigma

OpenStudy (amistre64):

thats the z score ... how do we use that to find probability?

OpenStudy (amistre64):

are you using a stats calculator like a ti83? or tables?

OpenStudy (skyepurple1997):

tables

OpenStudy (skyepurple1997):

the z score table

OpenStudy (amistre64):

is it a table from a book, or something else? there should be a *note on it for high z values.

OpenStudy (skyepurple1997):

my teacher gave me the table.... yeah, it says for z values greater than 3.49 use 1.000 to app. the area

OpenStudy (skyepurple1997):

is that what you're talking about? what does that mean?

OpenStudy (amistre64):

that tells you that for zscores larger than 3.49 the probability is 1 for all practical purposes what would you say that tells us about zscores less then -3.49?

OpenStudy (skyepurple1997):

on the negatiove side it says for values of z less than -3.49 use 0.000 to app. the area

OpenStudy (amistre64):

yeah ... so we should use 0.000 right?

OpenStudy (skyepurple1997):

i guess so...my teacher never talked about this. so, i have no idea

OpenStudy (amistre64):

is -3.9 less then -3.49?

OpenStudy (skyepurple1997):

yes

OpenStudy (skyepurple1997):

so, based on that statement, it would 0.000

OpenStudy (amistre64):

correct

OpenStudy (amistre64):

rnd to 4 decimals if your instructions are correct

OpenStudy (skyepurple1997):

so, 0.0000 lol

OpenStudy (amistre64):

better :)

OpenStudy (skyepurple1997):

:) thank you

OpenStudy (amistre64):

yw

OpenStudy (kvnmurty):

calculation of normal distribution variable Z = -3.928..... students t = -3.8888 with 49 degrees of freedom. If you look in the tables, the probability (one sided) for Z < -3.928 is 0... or for t < -3.888 it is 0. So answer is 0... It is correct.

OpenStudy (kvnmurty):

There is no need to be alarmed... As standard deviation of given distribution is very small. The standard error is very small: 0.128..for a mean of 88.. So a mean of 87.5 means dropping by 0.5 which is more than 3 times of standard error... So zero is good. Conceptually, probability of one truck load being < 87.5 is quite small, say "p". Now probability of each of 50 truck loads going below 87.5 simultaneously is p^50, which is almost 0.

OpenStudy (phi):

If you want the numerical value, wolfram can find it http://www.wolframalpha.com/input/?i=-3.928+standard+deviations

OpenStudy (phi):

it's pretty small z<-3.928 (left-tailed p-value) is 4.283×10^-5